OpenStudy (anonymous):

Kim is studying the sale of a particular brand of cereals from the year 2000 to 2010. She writes the following function to model the sale of the cereal S(t) in million dollars after t years: S(t) = t^2 + 7t + 69 What does the yintercept of the graph of the function represent. What is the reasonable domain. What is the average rate of change from 1 to 4.

OpenStudy (anonymous):

I REALLY REALLY NEED HELP ON THE DOMAIN I DID THE OTHER TWO

OpenStudy (anonymous):

@DecentNabeel

OpenStudy (decentnabeel):

\[\mathrm{Domain\:of\:}\:t^2+7t+69\::\quad \begin{bmatrix}\mathrm{Solution:}\: & \:-\infty \:<t<\infty \\\:\mathrm{Interval\:Notation:} & \:\left(-\infty \:,\:\infty \:\right)\end{bmatrix}\]

OpenStudy (anonymous):

How do you know

OpenStudy (anonymous):

So all real numbers

OpenStudy (decentnabeel):

yes

OpenStudy (anonymous):

Ok thanks so much

OpenStudy (decentnabeel):

wellcome :)

OpenStudy (anonymous):

@zepdrix